OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality 5x + y -10. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that includes the origin. Solid line with shading on the side that includes the origin. Dashed line with shading on the side that does not include the origin.

OpenStudy (anonymous):

Where is the inequality?

OpenStudy (anonymous):

its > with line underneath

OpenStudy (anonymous):

Assuming \(5x+y\ge-10\) Rearrange for y: \[y\ge-5x-10\] The 'line underneath' means that it is a solid line. So it is either A or C. This line, as it has as negative slope is slanted like '\', with the origin of the graph to the right. So, C. To check, the origin is (0, 0). Plug it in: \[5(0)+0\ge-10\] \[0\ge-10\] This is true. Therefore, it is correct.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

No problem.